OpenStudy (anonymous):

geometric mean of 3 and 300?

OpenStudy (aihberkhan):

What is \(3 * 300\)?

OpenStudy (anonymous):

no

OpenStudy (aihberkhan):

What? @clarkgrace

OpenStudy (aihberkhan):

I am asking you... what is 3 * 300?

OpenStudy (anonymous):

if b is geometric mean bean between a and c then b^2=ac \[b=\sqrt{ac}\]

OpenStudy (anonymous):

bean is typing error only.

OpenStudy (mathmale):

Please look up "geometric mean" on the 'Net before you try to calculate it. I found I had it all wrong and learned something valuable from this lookup.

OpenStudy (anonymous):

http://www.freemathhelp.com/geometric-mean.html

OpenStudy (anonymous):

multiply it together square root of 900

OpenStudy (anonymous):

A.B=C^2

OpenStudy (anonymous):

The formula to find the geometric mean is \[\sqrt[n]{a*b*c*d*...}\] Let a,b,c,d,... be the values given Let n be the number of values given Since the values given are 3 and 300, you just follow the formula \[\sqrt[2]{3*300}\] The answer would be 30 I hope this helped. :) Good luck!